5. Find the current is(t) using mesh analysis in the s-domain. (20 pts) F 4H mm HH 4e-2t u(t) V i2(1) 2eu(t) V 102
Answer #1

А L[a é at uts] = له. sta - induchor =.sl. Capacitor се 3 دراما 45 301 - 4 2 + low +) M S+2 I(3) IL (3) S+1 Loop I 4 + T (5)

2 Loop 2? AS[IG) - st 20 + 10[1411 – 5 (5) -10 11(1) + 1:09 [ 43+10] 2 9 +1 2 I() [43 +0] 371 +10 I, (5) S+) 2 10 IL (5) I(s)

3 20 I, (5) 3(43+1) + 10 s (45+1) +105 (5+1)(49+10) S(43+1) 20 125+3 +10[45+3)+103 I, (s) (3+1) (49+10) s(4+1) Il(s) 123 +3 +

20x3 (45+1) I, (3) (3+1)(43+0) 40 [s+ mesta 1.5 (48+1) 11() = (3+) (45+io) $ +0.83 + 0-0 -075) Bopay Rodos -0.870.8-4(2071) 2

Aa 2.7272 1. () +1) B - (⑨) (43+lo) (estio) (+) ( -10 4 +0.10345) o.3754 ST 10 뿜 B = ②.081 C= (④). (8+ 010845) S= -0.10845 1.

6 By substiteling A, B, CD 2.036028 -2.7272 2.001 0:17078 + Ir (9) + S+0.69154 3+1 3+0.10845 (43+10) -2.7272 2.081 0.17078 I,